Let å = (3. – 4) and 6 = ( - 1,5). Compute the dot product. å.b= Preview
Answer #1

Feel free to ask if you have any doubts and rate my answer with thumbs up if you like it!

We can weite 2 - 3 - 4 * 6=-1? + 59 (2.b) = (ax 6x) + (by ay) fani - RK 217 2.5 = (3x (-1)) + (-4xS) = -23) 2 - 3 - 20